Jose is twice as old as Stephanie, who is 7 years older than Andrew. Together their ages add to 41
Answer by ikleyn(52787)

.

Andrew + Stephanie + Jose    = 41


  x    +   (x+7)   + 2*(x+7) = 41


  x    +    x+7    + 2x + 14 = 41

  4x                         = 41 - 14 - 7

  4x                         = 20

   x                         = 20/4 = 5


You do the rest.

Solved.